OpenStudy (anonymous):

Solve. 3y + 15 > –6 (Points : 1) y > 3 y > –7 y > 13 y > –17

OpenStudy (anonymous):

3x=-6 -15=-21 x=-21/3 x=-7 then x>-7

OpenStudy (anonymous):

elimination method

OpenStudy (anonymous):

y>-7

OpenStudy (anonymous):

...its y and its a > or <

OpenStudy (anonymous):

its >